OpenStudy (anonymous):

The average of two numbers is 10.what is the sum of the two numbers

OpenStudy (alexandervonhumboldt2):

\[\frac{ a+b }{ 2 }=10\] multiply both sides by 2. what do you get?

OpenStudy (alexandervonhumboldt2):

\[\frac{ a+b }{ 2 }*2=10*2\] can you simplify this?